# Explicit Min-wise Hash Families with Optimal Size

### v1

@@ Abstract
We study explicit constructions of min-wise hash families and their extension to $k$-min-wise hash families. Informally, a min-wise hash family guarantees that for any fixed subset $X\subseteq[N]$, every element in $X$ has an equal chance to have the smallest value among all elements in $X$; a $k$-min-wise hash family guarantees this for every subset of size $k$ in $X$. Min-wise hash is widely used in many areas of computer science such as sketching [Coh16], web page detection [Hen06], and $\ell_{0}$ sampling [CF14]. For applications like similarity estimation [CDF+01] and rarity estimation [DM02], the space complexity of their streaming algorithms is roughly equal to the number of random bits used to construct such families.
The classical works by Indyk [Ind01] and Pătraşcu and Thorup [PT16] have shown $\Theta(\log(1/\delta))$-wise independent families give min-wise hash of multiplicative (relative) error $\delta$, resulting in a construction with $\Theta(\log(1/\delta)\log N)$ random bits. While this is optimal for constant errors, it leaves a gap to the existential bound of $O(\log(N/\delta))$ bits whenever $\delta$ is sub-constant, which is needed in several applications. Based on a reduction from pseudorandom generators for combinatorial rectangles by Saks, Srinivasan, Zhou and Zuckerman [SSZZ00], Gopolan and Yehudayoff [GY20] improved the number of bits to $O(\log N\log\log N)$ for polynomially small errors $\delta$. However, no construction with $O(\log N)$ bits (polynomial size family) and sub-constant error was known before.
@@ 1 Introduction
Min-wise hash families play a crucial role in the design of graph algorithms and streaming algorithms. Notable applications include similarity estimation [CDF+01], rarity estimation [DM02], data mining [HGKI02, Hen06], sketching [Coh16], and $\ell_{0}$ sampler [CF14, McG14]. In this work, we study explicit constructions of min-wise hash families with small size. In pseudorandomness, this is equivalent to studying the seed length (number of random bits used) to generate a hash function. In big data algorithms, this is the space complexity of applying min-wise hash families.
Following the standard notation of the min-wise hash [BCFM00, Ind01, FPS11], we consider multiplicative (relative) errors with respect to the fair probability in this work. Although this is different from the standard additive errors in pseudorandomness, multiplicative errors are crucial for many algorithmic applications of min-wise hash such as similarity estimation [CDF+01] and $\ell_{0}$ sampling [McG14].
@@
In this work, we will focus on the case of $M=\Omega(N/\delta)$ such that the uniform distribution over all functions from $[N]$ to $[M]$ satisfies (1.1) and (1.2) [Ind01, FPS11]. Specifically, let $U$ be the uniform distribution over all functions $h:[N]\to[M]$. Then $M=\Omega(N/\delta)$ implies
@@
This is equivalent to the requirement $|X|=O(\delta N)$ for $M=N$ in previous works [SSZZ00, Ind01, FPS11]. Moreover, most applications of min-wise hash could choose the image size $|M|$ to guarantee (1.3).
@@
Although ($k$-)min-wise hash has found a variety of applications in computer science, the primary approaches of explicit constructions are based on $t$-wise independent hash families [Ind01, FPS11] and pseudorandom generators for combinatorial rectangles [SSZZ00, GY20]. For min-wise hash, Indyk [Ind01] showed that any $O(\log(1/\delta))$-wise independent family is a min-wise hash family of error $\delta$. Since $t$-wise independent families need $\Theta(t\log NM)$ random bits, this construction needs $O(\log(1/\delta)\log NM)$ bits. In fact, Pătraşcu and Thorup [PT16] showed a matching lower bound: $\Omega(\log(1/\delta))$-wise independence is necessary to have error $\delta$. In contrast, non-explicitly it is known that one can use $O(\log(NM/\delta))$ random bits to construct min-wise hash families of error $\delta$. Therefore, although the construction in [Ind01] is optimal for constant errors, it fails to be optimal whenever the error is sub-constant.
@@
For $k$-min-wise hash, Feigenblat, Porat and Shiftan [FPS11] showed that $t$-wise independent family is also a $k$-min-wise hash family of error $\delta$ when $t=O(\log(1/\delta)+k\log\log(1/\delta))$. In turn, this construction needs $O\big((\log(1/\delta)+k\log\log(1/\delta))\cdot\log NM\big)$ random bits, which still leaves a gap to the optimal result of $O\big(k\log NM+\log(1/\delta)\big)$ bits for sub-constant $\delta=o(1)$.
@@
At the same time, Saks, Srinivasan, Zhou and Zuckerman [SSZZ00] reduced the construction of min-wise hash to pseudorandom generators for combinatorial rectangles of polynomially small errors. This reduction translates polynomially small (additive) errors to a multiplicative error like $\delta/|X|$ (described in Appendix A). Based on this reduction, Gopolan and Yehudayoff [GY20] provided a min-wise hash family of $O(\log NM\log\log NM)$ bits for any polynomially small error. Although this improves the result by Indyk [Ind01] of $O(\log^{2}NM)$ bits when $\delta$ is polynomially small, it does not provide a construction with $O(\log NM)$ bits even when $\delta$ is a constant.
@@
In this work, we study explicit constructions of min-wise hash with small sizes and (almost) polynomially small errors. Our constructions are well motivated, given that in practice, some applications of min-wise hash require small errors in which the seed length becomes the bottleneck on the space complexity of streaming algorithms. For example, a primary application of min-wise hash is $\ell_{0}$ sampling in the streaming model. Many graph streaming algorithms need $(k>1)$-min-wise hash with a sub-constant error $\delta=o(1)$ (see Table 1 in [KNP+17]) and have space complexity that is equal to the seed length of the min-wise hash family times the number of hashes used [AGM12]. Also, similarity estimation [CDF+01, DM02] applies $k$-min-wise hash of error $\delta$ directly to approximate the Jaccard similarity coefficient $\hat{S}(A,B):=\frac{|A\cap B|}{|A\cup B|}$ between two sets $A$ and $B$ as $(1\pm\delta)\cdot\frac{|A\cap B|}{|A\cup B|}\pm\frac{O(1)}{\delta\cdot k}$, so the space complexity is equal to the seed length of the $k$-min-wise hash family here.
@@
Furthermore, from a different aspect, given the connection between min-wise hash families and PRGs for combinatorial rectangles shown by Saks, Srinivasan, Zhou and Zuckerman [SSZZ00], a natural direction is to apply the results on the long line of research on PRGs for combinatorial rectangles to construct better min-wise hash families. Constructing pseudorandom generators for combinatorial rectangles have been extensively studied (e.g., [ASWZ96, LLSZ97, Lu02, GMR+12, GY20] to name a few) because they are related to fundamental problems in theoretical computer science such as derandomizing logspace computation and approximately counting the number of satisfying assignments of a CNF formula. While early works [ASWZ96, Lu02] in the 90s have already provided PRGs with seed length $O(\log NM)$ and slightly sub-constant errors (e.g., $2^{-\sqrt{\log NM}}$ [ASWZ96] and $2^{-\log^{2/3}NM}$ [Lu02]), no construction of min-wise hash family with $O(\log NM)$ bits and a sub-constant error was known before.
@@ 1.1 Our Results
We remark that the seed length of our min-wise hash is optimal up to constants and the error is almost polynomial up to a $\log\log N$ factor in the exponent. Hence Theorem˜1.2 improves previous results [SSZZ00, Ind01, FPS11, GY20] for seed length $O(\log N)$. In fact, this is the first construction of min-wise hash family with optimal seed length and sub-constant error.
@@
Again, this is the first construction of $k$-min-wise hash with optimal seed length and sub-constant error. One remark is that $k$-min-wise hash requires $\Omega(k\log N)$ bits even for a constant error. This is because the fair probability could be as small as $1/{N\choose k}$ in Definition (1.2). Also, as a direct application, our constructions give the optimal space complexity for many applications including similarity estimation and rarity estimation [CDF+01, DM02] whenever $k=\log^{O(1)}N$ and $\delta=2^{-O\left(\frac{\log N}{\log\log N}\right)}$.
@@ 1.2 Technique Overview
First of all, the connection shown in [SSZZ00] is not enough to directly use known PRGs for combinatorial rectangles to construct min-wise hash families with $O(\log N)$ bits. This is because one remarkable feature of min-wise hash is that the error is multiplicative with respect to the fair probability $1/|X|$, which could be as small as $1/N$. On the other hand, standard pseudorandom generators only consider additive errors, and constructing $O(\log NM)$ seed length PRGs fooling combinatorial rectangles with error $1/N$ is still a big open problem. More broadly, the long line of research on classical PRGs for read-once branching programs (ROBPs) ([Nis92, INW94, NZ96, BRRY10, BV10, KNP11, FK18, MRT19] to name a few) does not give a multiplicative error with $O(\log N)$ bits of seed.
@@
Since most PRGs for combinatorial rectangles and ROBPs are based on Nisan’s PRG [Nis92] (or its extension to the INW PRG [INW94]), a first attempt would be to modify these PRGs. However, the random event $(h(y)=\theta)$ already takes $\log_{2}M$ bits. Since $y$ could be any element, it is unclear how to revise these PRGs to replenish so many random bits just for one step (of $h(y)$).
@@
We consider a special type of extractors, which provides a strong guarantee when the source is uniform. Observe that it never hurts to put $h(y)$ as the first input of the ROBP under the Nisan-Zuckerman PRG. Suppose $y$ is in bucket $j\in[\ell]$ such that the value $h(y)$ is generated by $\mathsf{Ext}(w,s_{j})$. Our observation here is that as the first input, the random source $w$ is uniform at the beginning, hence one could expect stronger properties for $\mathsf{Ext}(w,s_{j})$ than for other $\mathsf{Ext}(w,s_{i})$ where $i\neq j$. In particular, we build an extractor such that $\mathsf{Ext}(U_{n},s)$ is uniform for a uniform source and any fixed seed. The construction is based on the sequence of works in pseudorandomness that designed linear seeded randomness extractors [NZ96, Tre01, GUV09].
@@
with a multiplicative error $\delta$. On the one hand, $O(1)$-wise independent family would guarantee $\Pr_{h\sim\mathcal{H}}[h(y)=\theta]=\Pr_{h\sim U}[h(y)=\theta]$. On the other hand, $\mathbf{1}(\min h(X\setminus y)>\theta)$ is a combinatorial rectangle where several PRGs [ASWZ96, Lu02, GMR+12, GY20] can fool it with a small additive error using $\widetilde{O}(\log N)$ bits of seed. This suggests us to consider the direct sum of a $t$-wise independent family and a PRG for combinatorial rectangles (see definition in Section˜2). However, it is unclear how to argue that this sum fools the conditional event $\Pr_{h\sim\mathcal{H}}[\min h(X\setminus y)>\theta\mid h(y)=\theta]$.
@@ 1.3 Discussion
    The fact that the Nisan-Zuckerman PRG can fool any input order has been used to fool formulas and general branching programs [IMZ19]. Can we find more applications of this powerful framework?
@@ 1.4 Related Works
Min-wise hash was introduced and investigated by Broder, Charikar, Frieze and Mitzenmacher [BCFM00] where the first definition set the probability to be exact $1/|X|$. This is equivalent to requiring that each function in the family is a permutation when $M=N$. Such a family is also called min-wise permutation. However, they showed a lower bound $\Omega(N)$ on the number of bits for the exact probability, and suggested to consider min-wise hash with multiplicative (relative) error for applications like similarity estimation and duplicate detection.
Later on, Indyk showed the first construction that $O(\log(1/\delta))$-wise independent families are min-wise hashing of error $\delta$. A matching lower bound on $t$-wise independent family was shown by Pătraşcu and Thorup [PT16] later.
For polynomially small errors, Saks, Srinivasan, Zhou and Zuckerman [SSZZ00] provided a construction of $O(\log^{3/2}N)$ bits based on the PRGs for combinatorial rectangles; this was improved to $O(\log N\log\log N)$ by Gopolan and Yehudayoff [GY20]. This work [GY20] is still the state-of-the-art of both PRGs for combinatorial rectangles and min-wise hash given polynomially small errors.
While one could run $k$ parallel min-wise hash to sample $k$ elements with replacement, $k$-min-wise looks for a sampling without replacement. This turns out to be more accurate in practice [CDF+01, DM02]. Feigenblat, Porat and Shiftan [FPS11] showed that $O(\log(1/\delta)+k\log\log(1/\delta))$-wise independent families are $k$-min-wise hashing of error $\delta$. Based on PRGs for combinatorial rectangles, Gopolan and Yehudayoff [GY20] constructed $k$-min-wise hash of $O\big(k\log N\cdot\log(k\log N)\big)$ bits for polynomially small errors.
Min-wise hash families and combinatorial rectangles are subclasses of read-once branching programs. So the classical PRG by Nisan [Nis92] of $O(\log^{2}N)$ bits implies a min-wise hash family of the same seed length. A long line of research has studied the effects and limitations of Nisan’s PRG (to name a few [INW94, BRRY10, BV10, De11, KNP11]). However, there has been little quantitative progress on the improvement of Nisan’s PRG. One exception is the construction of PRGs for combinatorial rectangles, where subsequent works [ASWZ96, Lu02] have reduced the seed length to $O(\log N)$ and achieved smaller errors.
The Nisan-Zuckerman PRG [NZ96] provides another method to derandomize ROBPs of seed length $O(\log N)$. While its output length is just $\log^{O(1)}N$, it can fool ROBPs of any input order. Impagliazzo, Meka and Zuckerman [IMZ19] have used this property to fool general formulas and branching programs.
Another powerful paradigm to design PRGs for ROBPs is via milder restrictions. Several beautiful applications are the PRGs for combinatorial rectangles and read-once CNFs [GMR+12, GKM15, GY20], the PRGs for ROBPs with an arbitrary input order [FK18], and the PRGs for ROBPs of width 3 [MRT19].

### v3

@@ Abstract
We study explicit constructions of min-wise hash families and their extension to $k$-min-wise hash families. Informally, a min-wise hash family guarantees that for any fixed subset $X\subseteq[N]$, every element in $X$ has an equal chance to have the smallest value among all elements in $X$; a $k$-min-wise hash family guarantees this for every subset of size $k$ in $X$. Min-wise hash is widely used in many areas of computer science such as sketching [Cohen2016], web page detection [Henzinger06], and $\ell_{0}$ sampling [sampling_survey]. For applications like similarity estimation [Cohen_similarity_est] and rarity estimation [DM_similarity_rarity], the space complexity of their streaming algorithms is roughly equal to the number of random bits used to construct such families.
The classical works by Indyk [minwise_Indyk] and Pătraşcu and Thorup [PatrascuT16] have shown $\Theta(\log(1/\delta))$-wise independent families give min-wise hash of multiplicative (relative) error $\delta$, resulting in a construction with $\Theta(\log(1/\delta)\log N)$ random bits. While this is optimal for constant errors, it leaves a gap to the existential bound of $O(\log(N/\delta))$ bits whenever $\delta$ is sub-constant, which is needed in several applications. Based on a reduction from pseudorandom generators for combinatorial rectangles by Saks, Srinivasan, Zhou and Zuckerman [SSZZ00], Gopalan and Yehudayoff [GY15] improved the number of bits to $O(\log N\log\log N)$ for polynomially small errors $\delta$. However, no construction with $O(\log N)$ bits (polynomial size family) and sub-constant error was known before.
@@ 1 Introduction
Min-wise hash families play a crucial role in the design of graph algorithms and streaming algorithms. Notable applications include similarity estimation [Cohen_similarity_est], rarity estimation [DM_similarity_rarity], data mining [HGKI02_data_mining, Henzinger06], sketching [Cohen2016], and $\ell_{0}$ sampler [sampling_survey, McGregor_survey]. In this work, we study explicit constructions of min-wise hash families with small size. In pseudorandomness, this is equivalent to studying the seed length (number of random bits used) to generate a hash function. In big data algorithms, this is the space complexity of applying min-wise hash families.
Following the standard notation of the min-wise hash [BRODER2000630, minwise_Indyk, minwise_FPS11], we consider multiplicative (relative) errors with respect to the fair probability in this work. Although this is different from the standard additive errors in pseudorandomness, multiplicative errors are crucial for many algorithmic applications of min-wise hash such as similarity estimation [Cohen_similarity_est] and $\ell_{0}$ sampling [McGregor_survey].
@@
In this work, we will focus on the case of $M=\Omega(N/\delta)$ such that the uniform distribution over all functions from $[N]$ to $[M]$ satisfies (1.1) and (1.2) [minwise_Indyk, minwise_FPS11]. Specifically, let $U$ be the uniform distribution over all functions $h:[N]\to[M]$. Then $M=\Omega(N/\delta)$ implies
@@
This is equivalent to the requirement $|X|=O(\delta N)$ for $M=N$ in previous works [SSZZ00, minwise_Indyk, minwise_FPS11]. Moreover, most applications of min-wise hash could choose the image size $|M|$ to guarantee (1.3).
@@
Although ($k$-)min-wise hash has found a variety of applications in computer science, the primary approaches of explicit constructions are based on $t$-wise independent hash families [minwise_Indyk, minwise_FPS11] and pseudorandom generators for combinatorial rectangles [SSZZ00, GY15]. For min-wise hash, Indyk [minwise_Indyk] showed that any $O(\log(1/\delta))$-wise independent family is a min-wise hash family of error $\delta$. Since $t$-wise independent families need $\Theta(t\log NM)$ random bits, this construction needs $O(\log(1/\delta)\log NM)$ bits. In fact, Pătraşcu and Thorup [PatrascuT16] showed a matching lower bound: $\Omega(\log(1/\delta))$-wise independence is necessary to have error $\delta$. In contrast, non-explicitly it is known that one can use $O(\log(NM/\delta))$ random bits to construct min-wise hash families of error $\delta$. Therefore, although the construction in [minwise_Indyk] is optimal for constant errors, it fails to be optimal whenever the error is sub-constant.
@@
For $k$-min-wise hash, Feigenblat, Porat and Shiftan [minwise_FPS11] showed that $t$-wise independent family is also a $k$-min-wise hash family of error $\delta$ when $t=O(\log(1/\delta)+k\log\log(1/\delta))$. In turn, this construction needs $O\big((\log(1/\delta)+k\log\log(1/\delta))\cdot\log NM\big)$ random bits, which still leaves a gap to the optimal result of $O\big(k\log NM+\log(1/\delta)\big)$ bits for sub-constant $\delta=o(1)$.
@@
At the same time, Saks, Srinivasan, Zhou and Zuckerman [SSZZ00] reduced the construction of min-wise hash to pseudorandom generators for combinatorial rectangles of polynomially small errors. This reduction translates polynomially small (additive) errors to a multiplicative error like $\delta/|X|$ (described in Appendix A). Based on this reduction, Gopalan and Yehudayoff [GY15] provided a min-wise hash family of $O(\log NM\log\log NM)$ bits for any polynomially small error. Although this improves the result by Indyk [minwise_Indyk] of $O(\log^{2}NM)$ bits when $\delta$ is polynomially small, it does not provide a construction with $O(\log NM)$ bits even when $\delta$ is a constant.
@@
In this work, we study explicit constructions of min-wise hash with small sizes and (almost) polynomially small errors. Our constructions are well motivated, given that in practice, some applications of min-wise hash require small errors in which the seed length becomes the bottleneck on the space complexity of streaming algorithms. For example, a primary application of min-wise hash is $\ell_{0}$ sampling in the streaming model. Many graph streaming algorithms need $(k>1)$-min-wise hash with a sub-constant error $\delta=o(1)$ (see Table 1 in [KNPWWY17]) and have space complexity that is equal to the seed length of the min-wise hash family times the number of hashes used [AGM12]. Also, similarity estimation [Cohen_similarity_est, DM_similarity_rarity] applies $k$-min-wise hash of error $\delta$ directly to approximate the Jaccard similarity coefficient $\hat{S}(A,B):=\frac{|A\cap B|}{|A\cup B|}$ between two sets $A$ and $B$ as $(1\pm\delta)\cdot\frac{|A\cap B|}{|A\cup B|}\pm\frac{O(1)}{\delta\cdot k}$, so the space complexity is equal to the seed length of the $k$-min-wise hash family here.
@@
Furthermore, from a different aspect, given the connection between min-wise hash families and PRGs for combinatorial rectangles shown by Saks, Srinivasan, Zhou and Zuckerman [SSZZ00], a natural direction is to apply the results on the long line of research on PRGs for combinatorial rectangles to construct better min-wise hash families. Constructing pseudorandom generators for combinatorial rectangles have been extensively studied (e.g., [ASWZ96_PRG_rect_comb, LLSZ95, Lu02, GMRTV, GY15] to name a few) because they are related to fundamental problems in theoretical computer science such as derandomizing logspace computation and approximately counting the number of satisfying assignments of a CNF formula. While early works [ASWZ96_PRG_rect_comb, Lu02] in the 90s have already provided PRGs with seed length $O(\log NM)$ and slightly sub-constant errors (e.g., $2^{-\sqrt{\log NM}}$ [ASWZ96_PRG_rect_comb] and $2^{-\log^{2/3}NM}$ [Lu02]), no construction of min-wise hash family with $O(\log NM)$ bits and a sub-constant error was known before.
@@ 1.1 Our Results
We remark that the seed length of our min-wise hash is optimal up to constants and the error is almost polynomial up to a $\log\log N$ factor in the exponent. Hence Theorem˜1.2 improves previous results [SSZZ00, minwise_Indyk, minwise_FPS11, GY15] for seed length $O(\log N)$. In fact, this is the first construction of min-wise hash family with optimal seed length and sub-constant error.
@@
Again, this is the first construction of $k$-min-wise hash with optimal seed length and sub-constant error. One remark is that $k$-min-wise hash requires $\Omega(k\log N)$ bits even for a constant error. This is because the fair probability could be as small as $1/{N\choose k}$ in Definition (1.2). Also, as a direct application, our constructions give the optimal space complexity for many applications including similarity estimation and rarity estimation [Cohen_similarity_est, DM_similarity_rarity] whenever $k=\log^{O(1)}N$ and $\delta=2^{-O\left(\frac{\log N}{\log\log N}\right)}$.
@@ 1.2 Technique Overview
First of all, the connection shown in [SSZZ00] is not enough to directly use known PRGs for combinatorial rectangles to construct min-wise hash families with $O(\log N)$ bits. This is because one remarkable feature of min-wise hash is that the error is multiplicative with respect to the fair probability $1/|X|$, which could be as small as $1/N$. On the other hand, standard pseudorandom generators only consider additive errors, and constructing $O(\log NM)$ seed length PRGs fooling combinatorial rectangles with error $1/N$ is still a big open problem. More broadly, the long line of research on classical PRGs for read-once branching programs (ROBPs) ([Nisan92, INW94, NZ96, BRRY, BV, KNP11, FK18, MRT19] to name a few) does not give a multiplicative error with $O(\log N)$ bits of seed.
@@
Since most PRGs for combinatorial rectangles and ROBPs are based on Nisan’s PRG [Nisan92] (or its extension to the INW PRG [INW94]), a first attempt would be to modify these PRGs. However, the random event $(h(y)=\theta)$ already takes $\log_{2}M$ bits. Since $y$ could be any element, it is unclear how to revise these PRGs to replenish so many random bits just for one step (of $h(y)$).
@@
We consider a special type of extractors, which provides a strong guarantee when the source is uniform. Observe that it never hurts to put $h(y)$ as the first input of the ROBP under the Nisan-Zuckerman PRG. Suppose $y$ is in bucket $j\in[\ell]$ such that the value $h(y)$ is generated by $\mathsf{Ext}(w,s_{j})$. Our observation here is that as the first input, the random source $w$ is uniform at the beginning, hence one could expect stronger properties for $\mathsf{Ext}(w,s_{j})$ than for other $\mathsf{Ext}(w,s_{i})$ where $i\neq j$. In particular, we build an extractor such that $\mathsf{Ext}(U_{n},s)$ is uniform for a uniform source and any fixed seed. The construction is based on the sequence of works in pseudorandomness that designed linear seeded randomness extractors [NZ96, Trevisan, GUV].
@@
with a multiplicative error $\delta$. On the one hand, $O(1)$-wise independent family would guarantee $\Pr_{h\sim\mathcal{H}}[h(y)=\theta]=\Pr_{h\sim U}[h(y)=\theta]$. On the other hand, $\mathbf{1}(\min h(X\setminus y)>\theta)$ is a combinatorial rectangle where several PRGs [ASWZ96_PRG_rect_comb, Lu02, GMRTV, GY15] can fool it with a small additive error using $\widetilde{O}(\log N)$ bits of seed. This suggests us to consider the direct sum of a $t$-wise independent family and a PRG for combinatorial rectangles (see definition in Section˜2). However, it is unclear how to argue that this sum fools the conditional event $\Pr_{h\sim\mathcal{H}}[\min h(X\setminus y)>\theta\mid h(y)=\theta]$.
@@ 1.3 Discussion
    The fact that the Nisan-Zuckerman PRG can fool any input order has been used to fool formulas and general branching programs [IMZ]. Can we find more applications of this powerful framework?
@@ 1.4 Related Works
Min-wise hash was introduced and investigated by Broder, Charikar, Frieze and Mitzenmacher [BRODER2000630] where the first definition set the probability to be exact $1/|X|$. This is equivalent to requiring that each function in the family is a permutation when $M=N$. Such a family is also called min-wise permutation. However, they showed a lower bound $\Omega(N)$ on the number of bits for the exact probability, and suggested to consider min-wise hash with multiplicative (relative) error for applications like similarity estimation and duplicate detection.
Later on, Indyk showed the first construction that $O(\log(1/\delta))$-wise independent families are min-wise hashing of error $\delta$. A matching lower bound on $t$-wise independent family was shown by Pătraşcu and Thorup [PatrascuT16] later.
For polynomially small errors, Saks, Srinivasan, Zhou and Zuckerman [SSZZ00] provided a construction of $O(\log^{3/2}N)$ bits based on the PRGs for combinatorial rectangles; this was improved to $O(\log N\log\log N)$ by Gopalan and Yehudayoff [GY15]. This work [GY15] is still the state-of-the-art of both PRGs for combinatorial rectangles and min-wise hash given polynomially small errors.
While one could run $k$ parallel min-wise hash to sample $k$ elements with replacement, $k$-min-wise looks for a sampling without replacement. This turns out to be more accurate in practice [Cohen_similarity_est, DM_similarity_rarity]. Feigenblat, Porat and Shiftan [minwise_FPS11] showed that $O(\log(1/\delta)+k\log\log(1/\delta))$-wise independent families are $k$-min-wise hashing of error $\delta$. Based on PRGs for combinatorial rectangles, Gopalan and Yehudayoff [GY15] constructed $k$-min-wise hash of $O\big(k\log N\cdot\log(k\log N)\big)$ bits for polynomially small errors.
Min-wise hash families and combinatorial rectangles are subclasses of read-once branching programs. So the classical PRG by Nisan [Nisan92] of $O(\log^{2}N)$ bits implies a min-wise hash family of the same seed length. A long line of research has studied the effects and limitations of Nisan’s PRG (to name a few [INW94, BRRY, BV, De, KNP11]). However, there has been little quantitative progress on the improvement of Nisan’s PRG. One exception is the construction of PRGs for combinatorial rectangles, where subsequent works [ASWZ96_PRG_rect_comb, Lu02] have reduced the seed length to $O(\log N)$ and achieved smaller errors.
The Nisan-Zuckerman PRG [NZ96] provides another method to derandomize ROBPs of seed length $O(\log N)$. While its output length is just $\log^{O(1)}N$, it can fool ROBPs of any input order. Impagliazzo, Meka and Zuckerman [IMZ] have used this property to fool general formulas and branching programs.
Another powerful paradigm to design PRGs for ROBPs is via milder restrictions. Several beautiful applications are the PRGs for combinatorial rectangles and read-once CNFs [GMRTV, GKM15, GY15], the PRGs for ROBPs with an arbitrary input order [FK18], and the PRGs for ROBPs of width 3 [MRT19].